OpenStudy (cluelessbookworm1042):

Use the slope formula to find the missing coordinate. m=-7/2, T(13,2), W(_, -5)

pooja195 (pooja195):

Hi there :) Welcome to Openstudy! :~) Here we use the slope formula which is: \[ \huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope} \] We are given several values so let's plug them in :) \[ \huge~\rm~\bf\color{blue}{ \frac{-5-2}{?-13}=\frac{-7}{2}} \]

pooja195 (pooja195):

We know -5-2=-7 So What minus 13 equals 2?

pooja195 (pooja195):

You can figure this out by doing the opposite of subtract, which is addition :) 13+2=?

OpenStudy (cluelessbookworm1042):

15

pooja195 (pooja195):

Good so out missing value is....?

OpenStudy (cluelessbookworm1042):

15

pooja195 (pooja195):

Good :)

OpenStudy (cluelessbookworm1042):

Thank you so much

pooja195 (pooja195):

You're welcome :-)